College of the Holy Cross MATH 135, section 1 Final Examination – Thursday, December 12

Your Name:

Instructions: For full credit, you must show all work on the test pages and place your final answer in the box provided for the problem. Use the back of the preceding page if you need more space for scratch work. The numbers next to each part of the questions are their point values.

Please do not write in the space below

Problem Points/Poss
I  / 25
II  / 20
III  / 20
IV  / 25
Total  /100